# Non-Linear Matter Power Spectrum Modeling in Interacting Dark Energy Cosmologies

### v1

@@ VI Final Remarks
In anticipation of future developments, our next steps involve the application and refinement of our models on real observational data, including datasets such as the Dark Energy Survey et al. (2022); Survey et al. (2023) and the Kilo-Degree Survey (KIDS-1000) Asgari et al. (2021b). Specifically, we aim to investigate the robustness of our model in addressing the current tension in the $S_{8}$ parameter. We are actively working on analyzing the results in this regard, and we look forward to sharing our findings in future communications.
@@ Acknowledgements.
We thank Carsten van de Bruck and Yuejia Zhai for useful discussions. E. S. thanks CAPES for partial financial support. R.C.N. thanks the financial support from the Conselho Nacional de Desenvolvimento Científico e Tecnologico (CNPq, National Council for Scientific and Technological Development) under the project No. 304306/2022-3, and the Fundação de Amparo à Pesquisa do Estado do RS (FAPERGS, Research Support Foundation of the State of RS) for partial financial support under the project No. 23/2551-0000848-3. E.D.V. is supported by a Royal Society Dorothy Hodgkin Research Fellowship. This article is based upon work from COST Action CA21136 Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by COST (European Cooperation in Science and Technology).

### v2

@@ VI Final Remarks
In anticipation of future developments, our next steps involve the application and refinement of our models on real observational data, including datasets such as the Dark Energy Survey et al. (2022); Abbott et al. (2023) and the Kilo-Degree Survey (KIDS-1000) Asgari et al. (2021b). Specifically, we aim to investigate the robustness of our model in addressing the current tension in the $S_{8}$ parameter. We are actively working on analyzing the results in this regard, and we look forward to sharing our findings in future communications.
@@ Acknowledgements.
The authors express their gratitude to the referee for their valuable comments and suggestions, which have significantly improved the quality of this work. We thank Carsten van de Bruck and Yuejia Zhai for useful discussions. E. S. thanks CAPES for partial financial support. R.C.N. thanks the financial support from the Conselho Nacional de Desenvolvimento Científico e Tecnologico (CNPq, National Council for Scientific and Technological Development) under the project No. 304306/2022-3, and the Fundação de Amparo à Pesquisa do Estado do RS (FAPERGS, Research Support Foundation of the State of RS) for partial financial support under the project No. 23/2551-0000848-3. E.D.V. is supported by a Royal Society Dorothy Hodgkin Research Fellowship. This article is based upon work from COST Action CA21136 Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by COST (European Cooperation in Science and Technology).